A 234.0 g piece of lead is heated to 83.0oC and then dropped into a calorimeter containing 585.0 g of water that initally is at 21.0oC. Neglecting the heat capacity of the container, find the final equilibrium temperature (in oC) of the lead and water.
Answer #1

Heat lost by lead block = heat gained by water

mlead * clead * (83 - T) = mwater * cwater * (T - 21)

Note that I have used 83 - T because there is a drop in temperature and T - 21 because there is a rise in temperature

234 g * 0.16 J / g oC  * ( 83 - T ) = 585 * 4.184 * (T - 21)

3107.52 - 37.44 * T = 2447.64 * T - 51400.44

54507.96 = 2485.08 * T

T = 21.93 oC